The table below offers EBIT for a potential capital investment for Fake Company Zeta. (This same project will be used for all of your FMC #3 work.) You should be able to determine a few things once you consider the following:

  • The initial investment is $20,000.
  • Depreciation is straight line over four years.
  • The company's WACC is estimated at 9.25%.
  • Company analysts estimate that a proper salvage value at the end of the project life of four years is 30% of the initial investment.
  • The company's tax rate is 30.0%.
YEAR 1 YEAR 2 YEAR 3 YEAR 4
EBIT $(1,750) $325 $1,025 $3,700

What is this project's net incremental cash flow in Year 2?
Answer #1

Cash flow year 2 = (EBIT)*(1-tax rate)+initial investment/depreciation period

=325*(1-0.3)+20000/4=5227.5